The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                 http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 57 July 21, 2001

Contents

Abstracts of 5 refereed papers  1
Abstracts of 3 non-refereed papers 4

Abstracts of Refereed Papers

900–1200Å interstellar extinction in the Galaxy, LMC, and SMC

J.B. Hutchings^{1} and J. Giasson^{1}

1 Herzberg Institute of Astrophysics, National Research Council of Canada, Victoria, B.C., Canada

We have derived far-ultraviolet extinction curves for stars in the Galaxy, LMC, and SMC, from spectra of pairs of stars observed with the FUSE satellite, matched in spectral type and with different amounts of reddening. The stars were observed for other FUSE programs and were not selected for large reddening. Thus, uncertainties in the derived extinction curves depend on the photometric data. In the case of strongly reddened stars from the Galaxy, the strong H_{2} absorption at the shortest wavelengths is an additional complication. These initial results indicate that there is a family of extinction curves that rise strongly to the Lyman limit, and extend the related extinction curves from longer wavelengths.

Accepted by: Publications of the Astronomical Society of the Pacific

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0107204

MSX, 2MASS, and the LMC: A Combined Near and Mid Infrared View

M.P. Egan^{1}, S.D. Van Dyk^{2}, and S.D. Price^{1}

1 Air Force Research Lab, USA 2 IPAC/Caltech, USA

The Large Magellanic Cloud (LMC) has been observed by both the Midcourse Space Experiment (MSX) in the mid-infrared and the Two Micron All Sky Survey (2MASS) in the near-infrared. We have performed a cross-correlation of the 1806 MSX catalog sources and nearly 1.4 million 2MASS catalogued point and extended sources and find 1664 matches. Using the available color information, we identify a number of stellar populations and nebulae, including main sequence stars, giant stars, red supergiants, carbon- and oxygen-rich asymptotic giant branch (AGB) stars, planetary nebulae, H II regions, and other dusty objects likely associated with early-type stars. 731 of these sources have no previous identification. We compile a listing of all objects, which includes photometry and astrometry. The 8.3 micron MSX sensitivity is the limiting factor for object detection: only the brighter red objects, specifically the red supergiants, AGB stars, planetary nebulae and H,H ii ii regions, are detected in the LMC. The remaining objects are likely in the Galactic foreground. The spatial distribution of the infrared LMC sources may contribute to understanding stellar formation and evolution and the overall galactic evolution. We demonstrate that a combined mid- and near-infrared photometric baseline provides a powerful means of identifying new objects in the LMC for future ground-based and space-based follow-up observations.

Accepted by: Astronomical Journal (2001 Oct issue)

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0107220

The Influence of Stellar Wind Variability on Measurements of Interstellar O vi Along Sightlines to Early-Type Stars

N. Lehner^{1}, A.W. Fullerton^{1,2}, K.R. Sembach^{1}, D.L. Massa^{3} and E.B. Jenkins^{4}

1 Department of Physics and Astronomy, Johns Hopkins University, 3400 N. Charles Street, Baltimore, MD 21218, USA 2 Department of Physics and Astronomy, University of Victoria, P.O. Box 3055, Victoria, BC V8W 3P6, Canada 3 Emergent IT, NASA’s Goddard Space Flight Center, Code 681, Greenbelt, MD 20771, USA 4 Princeton University Observatory, Princeton, NJ 08544, USA

A primary goal of the FUSE mission is to understand the origin of the O vi ion in the interstellar medium of the Galaxy and the Magellanic Clouds. Along sightlines to OB-type stars, these interstellar components are usually blended with O vi stellar wind profiles, which frequently vary in shape. In order to assess the effects of this time-dependent blending on measurements of the interstellar O vi lines, we have undertaken a mini-survey of repeated observations toward OB-type stars in the Galaxy and the Large Magellanic Cloud. These sparse time series, which consist of 2–3 observations separated by intervals ranging from a few days to several months, show that wind variability occurs commonly in O vi (∼60% of a sample of 50 stars), as indeed it does in other resonance lines. However, in the interstellar O vi λ1032 region, the O vi λ1038 wind varies only in ∼30% of the cases. By examining cases exhibiting large amplitude variations, we conclude that stellar-wind variability generally introduces negligible uncertainty for single interstellar O vi components along Galactic lines of sight, but can result in substantial errors in measurements of broader components or blends of components like those typically observed toward stars in the Large Magellanic Cloud. Due to possible contamination by discrete absorption components in the stellar O vi line, stars with terminal velocities greater than or equal to the doublet separation (1654 km s^{−1}) should be treated with care.

Accepted by: Astrophysical Journal Letter (astro-ph/0106556). For preprints, contact nl@pha.jhu.edu

The Thermal X-ray Emitting Shell of LMC Supernova Remnant 0549-69.3

Una Hwang^{1}, 2, Robert Petre^{1}, Stephen S. Holt^{3} and Andrew E. Szymkowiak^{4}

1 NASA Goddard Space Flight Center, Greenbelt, MD, USA 2 Department of Astronomy, University of Maryland, College Park, MD, USA 3 Olin College of Engineering, Needham, MA, USA

We use data from the Advanced CCD Imaging Spectrometer (ACIS) on the Chandra X-ray Observatory to image the shell surrounding the pulsar B0549-69, and to measure its spectral properties. Weak emission line features, notably of Fe, Mg and Si, indicate that the shell is composed primarily of ambient material heated by the blast wave. The shell emission in the east is faint and circular in appearance, with a temperature and ionization age of approximately 5 keV and 10^{10} cm^{−3} s, respectively. The emission in the west is brighter and has a more complex, distorted morphology, and is fitted with significantly lower average temperatures and higher ionization ages. An image with photon energies above 2 keV shows two hard arcs of emission located diametrically opposite each other at the outer boundary of the remnant shell; the spectra from these regions could include a substantial nonthermal component.

Accepted by: The Astrophysical Journal

Old Stellar Populations of the Small Magellanic Cloud

A. E. Dolphin^{1}, A. R. Walker^{2}, P. W. Hodge^{3}, M.Mateo^{4}, E. W. Olszewski^{5}, R. A. Schommer^{2}, N. B. Suntzeff^{2}

1Kitt Peak National Observatory, National Optical Astronomy Observatories, P.O. Box 26372, Tucson, AZ 85726, USA 2Cerro Tololo Inter-American Observatory, National Optical Astronomy Observatories, Casilla 603, La Serena, Chile 3Astronomy Department, University of Washington, Box 351580, Seattle, WA 98195-1580, USA 4Department of Astronomy, University of Michigan, Ann Arbor, MI 48109, USA 5Steward Observatory, University of Arizona, Tucson, AZ 85721, USA

We present WFPC2 and ground-based V I photometry of NGC 121 and a nearby field in the outer SMC. For NGC 121, we measure a true distance modulus of μ_{0} = 18.96 ± 0.04 (distance of 61.9 ± 1.1 kpc), age of 10.6±0.5 Gyr, metallicity of [Fe/H] = −1.03±0.06, and initial mass of 4.1±0.4×10^{5}M_{⊙}, assuming a Salpeter IMF with lower cutoff at 0.1M_{⊙}. In the outer SMC field, we find evidence of stars covering a wide range of ages – from 2 Gyr old to at least 9 − 12 Gyr old. We have measured the distance, extinction, and star formation history (past star formation rates and enrichment history) using a CMD-fitting algorithm. The distance modulus of the SMC is measured to be μ_{0} = 18.88±0.08, corresponding to a distance of 59.7 ± 2.2 kpc. The overall star formation rate appears to have been relatively constant over this period, although there may be small gaps in the star forming activity too small to be resolved. The lack of current star-forming activity is a selection effect, as the field was intentionally chosen to avoid recent activity. The mean metallicity of this field has increased from an average of [Fe/H] = −1.3 ± 0.3 for stars older than 8 Gyr to [Fe/H] = −0.7 ± 0.2 in the past 3 Gyr.

Accepted by: The Astrophysical Journal For preprints, contact dolphin@noao.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0107164

Abstracts of Non-Refereed Papers

The Shapes of Galaxies and Their Halos as Traced by Stars: The Milky Way Dark Halo and The LMC Disk

Roeland P. van der Marel^{1}

1 STScI, 3700 San Martin Drive, Baltimore, MD 21218, USA

Stars and their kinematics provide one of the tools available for studies of the shapes of galaxies and their halos. In this review I focus on two specific applications: the shape of the Milky Way dark halo and the shape of the LMC disk. The former is constrained by a variety of observations, but an accurate determination of the axial ratio q_{DH} remains elusive. A very flattened Milky Way dark halo with q_{DH} ≤ 0.4 is ruled out, and values q_{DH} ≥ 0.7 appear most consistent with the data. Near-IR surveys have revealed that the LMC disk is not approximately circular, as long believed, but instead has an axial ratio of 0.7 in the disk plane. The elongation is perpendicular to the Magellanic Stream, indicating that it is most likely due to the tidal force of the Milky Way. Equilibrium dynamical modelling of galaxies is important for many applications. At the same time, detailed studies of tidal effects and tidal streams have the potential to improve our understanding of both the Milky Way dark halo and the structure of satellite galaxies such as the LMC.

To appear in: Proceedings of the Yale Cosmology Workshop on ‘The Shapes of Galaxies and their Halos’, P. Natarajan, ed. (invited review) For preprints, contact marel@stsci.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0107248

T Tauri Stars in the Large Magellanic Cloud: a combined HST and VLT effort

M. Romaniello^{1}, N. Panagia^{2}, S. Scuderi^{3}, Roberto Gilmozzi^{4}, Eline Tolstoy^{5}, Fabio Favata^{6}, Robert P. Kirshner^{7}

1 European Southern Observatory – Garching bei München, Germany 2 Space Telescope Science Institute – Baltimore, USA 3 Osservatorio Astrofisico – Catania, Italy 4 European Southern Observatory – Paranal Observatory, Chile 5 Gemini Support Group – Oxford, United Kingdom 6 ESTEC – Noordwijk, The Netherlands 7 Harvard-Smithsonian Center for Astrophysics – Boston, USA

The combination of the unprecedented spatial resolution attainable with WFPC2 on board HST and of the large collecting area of the VLT makes it possible to study in detail the low mass pre-Main Sequence stars in galaxies other than our own. Here we present the results of our studies of two star forming environments in our closest galactic neighbor, the Large Magellanic Cloud: the region around Supernova 1987A and the double cluster NGC 1850.

To appear in: ”The Origins of Stars and Planets: the VLT View”, eds. João Alves and Mark McCaughrean For preprints, contact mromanie@eso.org

Also available from the URL http://arXiv.org/abs/astro-ph/0106321

Five Years of Magellanic Clouds Research: A Newsletter Editors’ Perspective

Eva K. Grebel^{1}, You-Hua Chu^{2}, John S. Gallagher^{3}, and Daniel Harbeck^{1}

1 Max Planck Institute for Astronomy, Königstuhl 17, D-69117 Heidelberg, Germany 2 Astronomy Department, University of Illinois, 1002 W. Green St., Urbana, IL 61801, USA 3 University of Wisconsin, Department of Astronomy, 475 N. Charter St., Madison, WI 53706, USA

We analyze the topical and demographic evolution of Magellanic Clouds research over the past five years based on submissions of abstracts of refereed papers to the electronic Magellanic Clouds Newsletter.

To appear in: “Dwarf Galaxies and their Environment”, Bad Honnef Workshop, January 2001, eds. K.S. de Boer, R.J. Dettmar, & U. Klein, Shaker Verlag

For preprints, contact grebel@mpia-hd.mpg.de
Also available from the URL http://arXiv.org/abs/astro-ph/0107143